OpenStudy (anonymous):

If (3*2^x)+(3*2^x)=128, what is the value of x?

OpenStudy (zzr0ck3r):

6*2^x=128 2^x=128/6 log(2^x)=log(128/6) x*log(2)=log(128/6) x=log(128/6) where log is log base 2

OpenStudy (zzr0ck3r):

log(128/6) = log(128)-log(6) = 7-log(6)

OpenStudy (zzr0ck3r):

again \[log(x) = \log_{2}(x) \]

OpenStudy (anonymous):

We know \(128=2^7\) because we're programmers (right? ;-) so we have:$$(3\cdot2^x)+(3\cdot2^x)=128\\3\cdot2^{x+1}=2^7\\3=2^{7-x-1}=2^{6-x}\\6-x=\log_23\\x=6-\log_23$$